# Imbalance-Robust and Sampling-Efficient Continuous Conditional GANs via Adaptive Vicinity and Auxiliary Regularization

### v2

@@ 4.2 Experimental Results
  • The two CcGAN-AVAR variants achieve superior performance across most settings by SFID metrics, maintaining comparable or better sample diversity than CcGAN/Dual-NDA while significantly improving label consistency - confirming the efficacy of their adaptive vicinity and auxiliary regularization designs. Crucially, CcGAN-AVAR preserves consistent generation quality between balanced (RC-49) and imbalanced (RC-49-I) datasets, unlike baseline methods that exhibit significant degradation, demonstrating exceptional robustness to data skews. Furthermore, the hybrid AV variant typically outperforms soft AV in label consistency across most configurations.
@@
Table 1: Comparison of generation quality and sampling efficiency at $64\times 64$ resolution. Performance is assessed using four metrics: SFID (overall quality), NIQE (visual realism), Diversity (sample diversity), and Label Score (label consistency), along with sampling speed and GPU memory consumption measured on an RTX 4090D. $\downarrow$” ($\uparrow$”) indicates lower (higher) values are preferred. The best and second-best results are highlighted in bold and underlined, respectively.
@@
Table 2: Comparison of generation quality and sampling efficiency at high resolutions ($64\times 64$) across methods. The best and second-best results are highlighted in bold and underlined, respectively.
@@ 4.3 Ablation Study
  • The Effects of AV and Regularization: Our component-wise analysis of CcGAN-AVAR on RC-49-I and UTKFace (Table 3) reveals that both soft and hybrid adaptive vicinity mechanisms substantially improve label consistency while maintaining or enhancing sample diversity. The auxiliary regularization terms further boost label consistency, resulting in significant overall quality improvements. It is worth noting that even the baseline configuration (without adaptive vicinity or auxiliary regularization) outperforms CcGAN/Dual-NDA (Table 1), due to a deeper GAN network and more advanced training techniques such as exponential moving average.
  • The Effects of $N_{\text{AV}}$: We investigate the impact of the key hyperparameter $N_{\text{AV}}$ on CcGAN-AVAR-H’s performance (Figure 7), conducting this ablation study with auxiliary regularization disabled to isolate AV’s effect. The results demonstrate substantial performance degradation when either disabling AV entirely ($N_{\text{AV}}=0$) or employing an excessively wide vicinity (large $N_{\text{AV}}$ values), while moderate $N_{\text{AV}}$ values consistently yield acceptable results. The dashed green lines in Figure 7 correspond to the optimal values selected for our main experiments.
@@
Soft AV Hybrid AV Aux. Reg. Aux. DRE
@@
Table 3: Ablation study: Component-wise analysis of CcGAN-AVAR
Figure 7: Ablation study: SFID/Label Score versus $N_{\text{AV}}$
5 Conclusion
This paper presents CcGAN-AVAR, an advanced CcGAN framework that achieves state-of-the-art generation performance while maintaining exceptional sampling efficiency. The proposed framework introduces two key innovations: (1) soft and hybrid adaptive vicinity mechanisms that effectively address the severe data imbalance inherent in regression datasets, replacing conventional fixed-size vicinity approaches; and (2) a dual regularization strategy incorporating both auxiliary regression and $f$-divergence penalties during generator training, significantly enhancing output quality. Extensive experiments across eight challenging imbalanced settings with varying resolutions demonstrate CcGAN-AVAR’s consistent superiority over existing methods.
6 Acknowledgments
This work was supported by the National Natural Science Foundation of China (Grant No. 62306147, 62201404, U24A20326), the Natural Science Foundation of Jiangsu Province (Grant No. BK20241901), Jiangsu Supplementary Funds of National Major Talent Project (2022-2024) (Grant No. R2024PT03), Nanjing Selective Funding Initiative for Returnee Innovation Projects (Grant No. R2024LZ02), the Startup Foundation for Introducing Talent of NUIST (Grant No. 2024r056), Yangtze River Delta Community of Sci-Tech Innovation (Grant No. 2023CSJZN0301), and Pujiang Talent Program (Grant No. 23PJ1412100).
@@ 6 Acknowledgments
Appendix A Adaptive Vicinity Implementation
A.1 Algorithm for Adaptive Vicinity Construction
Result: Vicinity width $\kappa_{y_{c}}$ and weight decay rate $v_{y_{c}}$
1 $\kappa_{l}\leftarrow 0$, $\kappa_{r}\leftarrow 0$;
2 $N_{c}=0$;
3 $t_{l}=i-1$, $t_{r}=i$;
// Vicinity extension
4 while $N_{c}<N_{\text{AV}}$ do
5 if $|y_{(t_{l})}-y_{c}|<|y_{(t_{r})}-y_{c}|$ then
    // Extend leftwards
6    $\kappa_{l}\leftarrow|y_{(t_{l})}-y_{c}|$;
7    $N_{c}\leftarrow N_{c}+N_{t_{l}}$;
8    $t_{l}\leftarrow t_{l}-1$;
9
10 else if $|y_{(t_{l})}-y_{c}|>|y_{(t_{r})}-y_{c}|$ then
    // Extend rightwards
11    $\kappa_{r}\leftarrow|y_{(t_{r})}-y_{c}|$;
12    $N_{c}\leftarrow N_{c}+N_{t_{r}}$;
13    $t_{r}\leftarrow t_{r}+1$;
14
15 else
    // Extend in both directions
16    $\kappa_{l}\leftarrow|y_{(t_{l})}-y_{c}|$;
17    $\kappa_{r}\leftarrow|y_{(t_{r})}-y_{c}|$;
18    $N_{c}\leftarrow N_{c}+N_{t_{l}}+N_{t_{r}}$;
19    $t_{l}\leftarrow t_{l}-1$;
20    $t_{r}\leftarrow t_{r}+1$;
21
22   end if
23
24 end while
[…]
26 $\nu_{y_{c}}\leftarrow 1/\kappa_{y_{c}}$;
A.2 Heuristic Selection of $N_{c}$
The minimum number of effective samples per vicinity ($N_{c}$) serves as the crucial hyperparameter in CcGAN-AVAR, with ablation studies on RC-49-I and UTKFace demonstrating optimal performance when $N_{c}$ is properly calibrated within a moderate range, as evidenced in Figure S.8.
To determine $N_{c}$ in practice, we propose a heuristic approach:
  • First, identify all the distinct training labels
    $Y^{u}=\{y_{(1)},...,y_{(M)}\}$ (S.11)
    and their sample counts
    $\mathcal{N}=\{N_{1},...,N_{M}\}.$ (S.12)
    $\mathcal{S}=\{S_{1},...,S_{M-1}\}.$ (S.13)
  • Finally, compute the averaged sample count $\bar{S}=\frac{1}{M-1}\sum_{i=1}^{M-1}S_{i}$, and set $N_{c}$ to an integer near $\bar{S}$ with potential minor adjustments based on the sparsity characteristics of the data distribution.
For instance, on the bimodal RC-49-I dataset where $\bar{S}\approx 25.279$, we set $N_{c}=30$. For UTKFace with $\bar{S}\approx 480$, we chose $N_{c}=400$. In the case of Steering Angle, where many $S_{i}\in\mathcal{S}$ cluster around 2 despite $\bar{S}\approx 13$, we selected $N_{c}=20$ to mitigate the impact of sparse label regions.
Figure S.8: Ablation analysis of adaptive vicinity size ($N_{c}$) effects on CcGAN-AVAR-H performance, comparing SFID and Label Score metrics across RC-49-I and UTKFace datasets at 64×64 resolution. Dashed green lines correspond to the parameter values used in the primary experiment. Auxiliary regularization was disabled to isolate the contribution of adaptive vicinity.
Appendix B Auxiliary Regularization Implementation
B.1 Network Architecture
CcGAN-AVAR adopts a SNGAN-based architecture (Miyato et al. 2018), incorporating an additional residual block in the generator compared to the original implementation in (Ding et al. 2023b). While Ding et al. (2023b) recommends SAGAN (Zhang et al. 2019) for high-resolution scenarios, we observe that its self-attention mechanism exhibits limited compatibility with exponential moving average (EMA) (Karras et al. 2024) – a technique that effectively improves SNGAN performance in our experiments.
A key contribution of this work is our modified discriminator architecture, which integrates both a two-layer regression branch and a three-layer density ratio estimation (DRE) branch, as illustrated in Figures S.9 and S.10. The regression branch employs two linear layers with spectral normalization (SN) (Miyato et al. 2018) and ReLU activation, while the DRE branch replaces SN with group normalization (GN) (Wu and He 2020) using 8 groups. For high-resolution UTKFace (192×192) experiments, we incorporate a dropout layer after ReLU activation - crucially, not for the conventional purpose of preventing overfitting, but to address a distinct training challenge where the DRE branch otherwise fails to converge (evidenced by stagnant training loss).
Complete architectural details and implementations are provided in our code repository.
Figure S.9: Architecture of the two-layer fully connected regression branch $f_{\text{reg}}$ in the discriminator.
Figure S.10: Architecture of the three-layer fully connected DRE branch $f_{\text{dre}}$ in the discriminator, where Dropout is optional.
B.2 Relation to ACGAN
CcGAN-AVAR differs fundamentally from ACGAN (Odena, Olah, and Shlens 2017) in three key aspects. First, while ACGAN uses an auxiliary classification branch for class-conditional generation, its discriminator does not directly receive class information. In contrast, CcGAN-AVAR integrates a regression branch for continuous label prediction and explicitly injects condition labels into the discriminator via label projection (Figure 3). Second, whereas ACGAN relies on a cross-entropy loss for classification, CcGAN-AVAR employs a regression loss—specifically, a custom $\gamma$-insensitive hinge loss (Eq. (6)) designed to improve robustness against label noise, rather than a standard mean squared error. Third, and most distinctively, CcGAN-AVAR implements a density ratio estimation branch that imposes an $f$-divergence penalty during generator training - a feature entirely absent in the standard ACGAN architecture.
B.3 Extra Illustration for $f$-Divergence Penalty
The $f$-divergence (Nowozin, Cseke, and Tomioka 2016) between $p_{r}(\bm{x}|y)$ and $p_{g}(\bm{x}|y)$ is defined as:
$=\mathbb{E}_{\bm{x}\sim p_{g}(\bm{x}|y)}\left[f\left(\frac{p_{r}(\bm{x}|y)}{p_{g}(\bm{x}|y)}\right)\right],$ (S.14)
$=\mathbb{E}_{\bm{x}\sim p_{g}(\bm{x}|y)}\left[(r(\bm{x}|y)-1)^{2}\right].$ (S.15)
The proposed $f$-divergence penalty of CcGAN-AVAR is defined as follows:
$\mathcal{L}^{G}_{f}=\frac{1}{N^{g}}\sum_{i=1}^{N^{g}}\left(f_{\text{dre}}(D_{\text{conv}}(\bm{x}_{i}^{g},y_{i}^{g}))-1\right)^{2}.$ (S.16)
In Eq. (S.16), $f_{\text{dre}}(D_{\text{conv}}(\bm{x}_{i}^{g},y_{i}^{g}))$ estimates the conditional density ratio $r(\bm{x}|y)$. Thus, Eq. (S.16) approximates Eq. (S.15), and minimizing $\mathcal{L}^{G}_{f}$ encourages the generator to better match the target data distribution.
Appendix C Experimental Setup and Implementation
C.1 How RC-49-I is made?
Since RC-49 is a relatively simple synthetic dataset with a sufficiently large sample size for reliable evaluation, we construct three imbalanced subsets to thoroughly assess the robustness of CcGAN-AVAR and other baseline methods. As shown in Figure S.11, these subsets exhibit unimodal, bimodal, and trimodal label distributions. Using the unimodal case as an example, we describe the derivation procedure from RC-49 as follows:
@@ C.1 How RC-49-I is made?
C.2 Hardware and Software Details
@@
  • $64\times 64$ resolution: 24GB GPU memory
  • $128\times 128$ and $192\times 192$ resolutions: 48GB GPU memory (RTX 4090D-48G)
@@
C.3 Training Configuration Details
@@
Table S.4 details our CcGAN-AVAR training configurations, which employ a modified SNGAN backbone with exponential moving average and mixed-precision training - improvements over Ding et al. (2023b) [^1] https://github.com/UBCDingXin/improved˙CcGAN and Ding, Wang, and Xu (2024) [^2] https://github.com/UBCDingXin/Dual-NDA. Through mixed-precision training, our implementation demonstrates a 2× improvement in training speed and 50% reduction in memory usage compared to Ding et al. (2023b) when using identical architecture and batch size configurations.
@@
C.4 Evaluation Configuration Details
Our evaluation protocol follows Ding et al. (2023b); Ding, Wang, and Xu (2024); Ding et al. (2025), with detailed specifications provided in Appendix S.V.B of Ding et al. (2025). For the RC-49-I experiments, we maintain identical evaluation procedures to those used for RC-49.
Appendix D Extra Experimental Results
Example synthetic images generated by Dual-NDA and CcGAN-AVAR for the UTKFace and Steering Angle experiments ($\times128$ resolution) are presented in Figs. S.12 and S.13.
Figure S.12: Some example fake images from Dual-NDA and CcGAN-AVAR-H in the UTKFace ($\times128$) experiment.
Figure S.13: Some example fake images from Dual-NDA and CcGAN-AVAR-H in the Steering Angle ($\times128$) experiment.

### v3

@@ IV-B Experimental Results
  • The two CcGAN-AVAR variants achieve superior performance across most settings by SFID metrics, maintaining comparable or better sample diversity than CcGAN/Dual-NDA while significantly improving label consistency—confirming the efficacy of their adaptive vicinity and auxiliary regularization designs. Crucially, CcGAN-AVAR preserves consistent generation quality between balanced (RC-49) and imbalanced (RC-49-I) datasets, unlike baseline methods that exhibit significant degradation, demonstrating exceptional robustness to data skews. Furthermore, the hybrid AV variant typically outperforms soft AV in label consistency across most configurations.
@@
  • The performance gap between CcGAN-AVAR and CCDM—in terms of both generation quality and sampling efficiency—widens significantly at higher resolutions, most notably at $256\times 256$. This is clearly demonstrated in the Steering Angle ($256\times 256$) experiment: CCDM’s anomalously high Label Score signifies a fundamental failure, while the CcGAN-AVAR variants achieve superior performance, attaining the lowest SFID and Label Scores alongside the highest sampling speed and lowest memory consumption.
TABLE I: Comparison of generation quality and sampling efficiency at $64\times 64$ resolution
@@
  • Generation quality is evaluated by SFID (overall quality), NIQE (visual realism), Diversity (sample diversity), and Label Score (label consistency). Sampling efficiency is measured by inference speed and GPU memory consumption on a single NVIDIA RTX 4090D-24G GPU. $\downarrow$” ($\uparrow$”) indicates lower (higher) values are preferred. The best and second-best results are highlighted in bold and underlined, respectively.
TABLE II: Comparison of generation quality and sampling efficiency at high resolutions ($64\times 64$) across methods
@@
UTKFace
($\times256$)
CcDPM (AAAI’24) 0.559 1.552 0.854 12.121 0.30 18.17
CCDM (arXiv’25) 0.268 1.364 1.121 7.555 0.30 23.51
CcGAN (T-PAMI’23) 0.347 1.308 1.053 6.010 384.62 12.50
Dual-NDA (AAAI’24) 0.319 1.381 1.216 5.712 384.62 12.50
CcGAN-AVAR-S 0.196 1.601 1.228 5.678 624.87 16.39
CcGAN-AVAR-H 0.196 1.347 1.219 5.798 624.87 16.39
Steering Angle
($\times256$)
CcDPM (AAAI’24) 1.217 1.748 1.330 29.239 0.30 18.17
CCDM (arXiv’25) 0.902 1.616 1.151 23.057 0.30 23.31
CcGAN (T-PAMI’23) 0.984 1.999 1.054 8.399 613.50 12.40
Dual-NDA (AAAI’24) 0.967 1.834 0.951 8.338 613.50 12.40
CcGAN-AVAR-S 0.718 2.177 1.111 6.522 626.17 11.97
CcGAN-AVAR-H 0.683 1.966 1.113 4.958 626.17 11.97
  • Generation quality is evaluated by SFID (overall quality), NIQE (visual realism), Diversity (sample diversity), and Label Score (label consistency). Sampling efficiency is measured by inference speed and GPU memory consumption on a single NVIDIA RTX 4090D-48G GPU. $\downarrow$” ($\uparrow$”) indicates lower (higher) values are preferred. The best and second-best results are highlighted in bold and underlined, respectively.
@@ IV-C Ablation Study
  • The Effects of AV and Regularization: Our component-wise analysis of CcGAN-AVAR on RC-49-I and UTKFace (Table III) reveals that both Soft and Hybrid Adaptive Vicinity (SAV/HAV) mechanisms substantially improve label consistency while maintaining or enhancing sample diversity. The auxiliary regularization terms further boost label consistency, resulting in significant overall quality improvements. It is worth noting that even the baseline configuration (without adaptive vicinity or auxiliary regularization) outperforms CcGAN/Dual-NDA (Table I), due to a deeper GAN network and more advanced training techniques such as exponential moving average.
    We also evaluate the effect of HAV by comparing the strong CcGAN baseline with and without it on the UTKFace dataset ($64\times 64$), as shown in Fig. 8. In this ablation study, auxiliary generator regularization was disabled to isolate HAV’s contribution. The results demonstrate that HAV consistently improves both sample diversity and label consistency. The bar chart at the bottom displays the sample size for each age.
  • The Effects of $N_{\text{AV}}$: We investigate the impact of the key hyperparameter $N_{\text{AV}}$ on CcGAN-AVAR’s performance (Fig. 9), conducting this ablation study with auxiliary regularization disabled to isolate AV’s effect. The results demonstrate substantial performance degradation when either disabling AV entirely ($N_{\text{AV}}=0$) or employing an excessively wide vicinity (large $N_{\text{AV}}$ values), while moderate $N_{\text{AV}}$ values consistently yield acceptable results. The dashed green lines in Fig. 9 correspond to the optimal values selected for our main experiments.
  • The Effect of Network Architecture: We further evaluate the generation performance of CcGAN-AVAR-H using different network architectures, including DCGAN [25], SNGAN [24], SAGAN [26], and BigGAN (including BigGAN-deep) [27]. The quantitative results, summarized in Table IV, show that the SNGAN backbone performs best on CCGM benchmark datasets. This outcome validates the network design we proposed in Section III-D.
TABLE III: Ablation study: Component-wise analysis of CcGAN-AVAR.
@@
Strong
CcGAN
Soft
AV
Hybrid
AV
Regression
Penalty
$f$-divergence
Penalty
@@
  • The quantitative results marked with an asterisk (*) correspond to the vanilla CcGAN implementation from [16, 14]. Our strong CcGAN baseline, implemented using the unified codebase with SNGAN backbone, exponential moving average (EMA), and mixed-precision training, achieves comparable or superior generation performance to the vanilla CcGAN while requiring substantially less training cost. By incrementally incorporating the proposed soft/hybrid adaptive vicinity, regression penalty, and $f$-divergence penalty into this baseline, the generation performance is progressively improved.
Figure 8: Performance of the strong CcGAN baseline with and without the proposed Hybrid Adaptive Vicinity (HAV) on UTKFace ($64\times 64$), with auxiliary generator regularization disabled. Sample sizes per age are indicated by the blue bar chart at the bottom in each image. HAV significantly improves overall sample diversity and label consistency, despite minor diversity reductions at a few specific ages.
Figure 9: Ablation analysis of adaptive vicinity size ($N_{\text{AV}}$) effects on CcGAN-AVAR-H performance, comparing SFID and Label Score metrics across RC-49-I and UTKFace datasets at 64×64 resolution. Dashed green lines correspond to the parameter values used in the primary experiment. Auxiliary regularization was disabled to isolate the contribution of adaptive vicinity.
TABLE IV: Ablation study: The impact of network architecture on CcGAN-AVAR-H.
Dataset
Network
Architecture
SFID
NIQE Diversity
Label
Score
RC-49-I
($64\times 64$)
SNGAN 0.050 1.729 3.706 1.402
DCGAN 0.063 1.737 3.636 1.585
SAGAN 0.082 1.972 3.631 1.360
BigGAN 0.071 1.958 3.650 1.458
UTKFace
($64\times 64$)
SNGAN 0.356 1.691 1.278 6.696
DCGAN 0.363 1.690 1.262 6.701
SAGAN 0.396 1.657 1.228 7.159
BigGAN 0.369 1.741 1.261 7.241
UTKFace
($128\times 128$)
SNGAN 0.297 1.173 1.251 6.586
SAGAN 0.378 1.255 1.183 6.759
BigGAN 0.337 1.130 1.215 7.464
BigGAN-deep 0.464 1.248 1.173 4.837
  • Since more complex architectures like SAGAN, BigGAN, and BigGAN-deep underperform SNGAN on the CCGM benchmarks, we adopt SNGAN as the backbone for our CcGAN-AVAR implementation.
V Conclusion
This paper presents CcGAN-AVAR, an advanced CcGAN framework that achieves state-of-the-art generation performance while maintaining exceptional sampling efficiency. The proposed framework introduces two key innovations: (1) soft and hybrid adaptive vicinity mechanisms that effectively address the severe data imbalance inherent in regression datasets, replacing conventional fixed-size vicinity approaches; and (2) a dual regularization strategy incorporating both auxiliary regression and $f$-divergence penalties during generator training, significantly enhancing output quality. Extensive experiments across eleven challenging imbalanced settings with varying resolutions demonstrate CcGAN-AVAR’s consistent superiority over existing methods.
@@ V Conclusion
Appendix A GitHub repository
The source code and implementation details will be made publicly available at:
https://github.com/UBCDingXin/CcGAN-AVAR
Appendix B Experimental Setup and Implementation
B-A How RC-49-I is made?
Since RC-49 is a relatively simple synthetic dataset with a sufficiently large sample size for reliable evaluation, we construct three imbalanced subsets to thoroughly assess the robustness of CcGAN-AVAR and other baseline methods. As shown in Fig. S.10, these subsets exhibit unimodal, bimodal, and trimodal label distributions. Using the unimodal case as an example, we describe the derivation procedure from RC-49 as follows:
@@ B-A How RC-49-I is made?
B-B How $256\times 256$ versions of UTKFace and Steering Angle are made?
UTKFace. We perform super-resolution on the $192\times 192$ version of UTKFace using pre-trained Real-ESRGAN [40] and GFPGAN [41]. First, Real-ESRGAN [^1] https://github.com/xinntao/Real-ESRGAN/releases/download/v0.1.0/RealESRGAN_x4plus.pth is employed to reconstruct $256\times 256$ images from the $192\times 192$ inputs, primarily enhancing background textures and overall image quality. Subsequently, GFPGAN [^2] https://github.com/TencentARC/GFPGAN/releases/download/v1.3.0/GFPGANv1.3.pth detects facial regions in the super-resolved images. After alignment, these regions are processed by a face restoration network. The restoration results are then inversely transformed to their original positions via an affine transformation and blended with the background using mask-based weighted fusion, with boundary feathering applied to minimize artifacts. Finally, the enhanced image is resampled to the target resolution and saved in a new HDF5 file, preserving the original data format.
Steering Angle. We apply the same pre-trained Real-ESRGAN to perform super-resolution on the $128\times 128$ Steering Angle dataset, saving the reconstructed images to a new HDF5 file.
Fig. S.11 shows some example images from both datasets before and after super-resolution. The datasets are available for download via the links provided in our code repository.
Figure S.11: Example images of UTKFace and Steering Angle before and after super-resolution.
B-C Hardware and Software Details
@@
  • $64\times 64$ resolution: RTX 4090D-24G
  • $64\times 64$ resolutions: RTX 4090D-48G (modified)
@@
B-D Training Configuration Details
@@
Table S.5 details our CcGAN-AVAR training configurations, which employ a modified SNGAN backbone with exponential moving average and mixed-precision training—improvements over [16] and [23]. Through mixed-precision training, our implementation demonstrates a 2× improvement in training speed and 50% reduction in memory usage compared to [16] when using identical architecture and batch size configurations.
@@
UTKFace
($\times256$)
SNGAN, gene_ch=64, disc_ch=48, hinge loss, DiffAugment, use EMA, steps=100K, batch size=64,
[…]
Steering Angle
($\times256$)
SNGAN, gene_ch=64, disc_ch=48, hinge loss, DiffAugment, use EMA, steps=80K (hybrid) / 30K (soft),
[…]
$\lambda_{\text{reg}}^{D}=\lambda_{\text{reg}}^{G}=1$, $\lambda_{\text{dre}}^{G}=0.5$, $\lambda_{f}^{G}=0.5$
@@
B-E Evaluation Configuration Details
Our evaluation protocol follows [16, 23, 14], with detailed specifications provided in Appendix S.V.B of [14]. For the RC-49-I experiments, we maintain identical evaluation procedures to those used for RC-49.
The sampling batch size was set to 200 for resolutions below $256\times 256$ and 100 for $256\times 256$ experiments.
Appendix C Extra Experimental Results
Example synthetic images generated by CcGAN-AVAR-H for the UTKFace and Steering Angle experiments ($\times256$ resolution) are presented in Figs. S.12 and S.13.
Figure S.12: Fake images generated by CcGAN-AVAR-H for the UTKFace dataset at $\times256$ resolution. Each row corresponds to a specific age.
Figure S.13: Fake images generated by CcGAN-AVAR-H for the Steering Angle dataset at $\times256$ resolution. Each row corresponds to a specific angle.